OpenStudy (anonymous):

PLEASE HELP!!! (SIMPLIFY EXPRESSIONS) i need work to go with it :/ fan + medal awarded :)) 7 + 5c + 8c^2 - 4c - 10 + 5c^2

Vocaloid (vocaloid):

start by combining like terms

Vocaloid (vocaloid):

8c^2 + 5c^2 = ?

OpenStudy (anonymous):

13^2 ?

Vocaloid (vocaloid):

don't forget the c

OpenStudy (anonymous):

13c^2

Vocaloid (vocaloid):

good. 5c - 4c = ?

OpenStudy (anonymous):

1c

Vocaloid (vocaloid):

good, but we usually write 1c as c by itself 7 - 10 = ?

OpenStudy (anonymous):

-3

Vocaloid (vocaloid):

good, now we just put everything together to get 13c^2 + c - 3

Vocaloid (vocaloid):

and that's it

OpenStudy (anonymous):

so that's the answer?

OpenStudy (anonymous):

thank you :3